y1=13x-4 and y2=5x+10 y1 exceeds y2 by 2
can you please help me set the problem up to solve for all variables of x?

Answer by ankor@dixie-net.com(22740)

y1 = 13x - 4 and y2 = 5x + 10. y1 exceeds y2 by 2
can you please help me set the problem up to solve for all variables of x?
:
y1 = (13x - 4)
:
y2 = (5x + 10)
:
y1 = y2 + 2; same as "y1 exceeds y2 by 2"
:
Substitute (13x - 4) for y1 in the above equation:
(13x - 4) = y2 + 2
:
Substitute (5x + 10) for y2 in the above equation:
13x - 4 = (5x + 10) + 2
:
13x - 4 = 5x + 12
:
13x - 5x = 12 + 4
:
8x = 16
:
x = 16/8
:
x = 2
:
:
Check solution by substitution in the original equations:
y1 = 13(2) - 4
y1 = 26 - 4
y1 = 22
and
y2 = 5(2) + 10
y2 = 10 + 10
y2 = 20
:
y1 does indeed exceed y2 by 2; confirms our solution
:
:
Did this help you understand this?